Logbooks general detector daq Analysis Physics 12 GeV HallA OldDVCS2 DVCS2Ana Parity
runliste issues procurement survey minutes web compton2 NewIssues NewProcurement
  Minutes  Not logged in ELOG logo
Message ID: 12     Entry time: 08/05/04, 16:03
Author: Franck Sabatie 
Subject: DVCS meeting minutes - 08/04/04 
  Present: Charles, Gagik, David, Catherine, Alex, Jean-Marc, Franck,
Carlos, Matthieu, Eric, Malek, Sirish.

  Franck/Installation:

  - Electronics rack will move into the Hall on thursday afternoon.
Calorimeter unstacking will take place in the clean room on
thursday-friday. Blocks will be brought 5 by 5 to the Hall for
repair/stocking/restacking. DVCS detectors+platform will go into the Hall
early next week. RG213 will be moved at about the same time.

  - Few remaining question to Jack: RG213 cable management (carriages?).

  Sirish/Documents:

  - Need to write COO, ESSAD, RSAD. TSOP may not be needed (will know
soon). ESSAD is what we should focus one. Each subsystem should write
about 1-2 pages describing installation and associated hazard and security
features.

  - Still some money to spend quickly.

  Gagik/Proton Array:

  - repaired one of the worst-case block (horizontal tower). Only took
20min to replace the base which didn't work. Now, all PA channels are
fully operational.

  Carlos/Calorimeter:

  - Performed tests using continuous LED. Gain stability was better than
1% at 20uA and better at lower anode current.

  - Performed stability tests, using chronical LED scans over a period of
10hrs. Clearly the system needs about 1-2 hrs to warm up. After this,
stability is better than 0.3% (might be even better, since the test is
limited by the statistical accuracy of the LED scan itself).

  Eric/Veto:

  - Patch panels are being repaired by local electronics group. Everything
will be ready by the end of the week.

  - Writing 2 tech. notes, will be available shortly.

  Alex/DAQ:

  - problem with 2 ARS boards. Had to use spare. Now no spare anymore.
Faulty boards will be repaired/changed. Monday-Wednesday next week, Alex
and Bob will work together on DAQ, and integration into Hall A.

  - Will try new PowerPC CPUs.

  - Some remote switched have been ordered. Might need one more.

  Franck/DB:

  - script to send coda information to MySQL database on a run by run
basis has been written.
ELOG V3.1.4-unknown